OpenStudy (anonymous):

x=a, y=b or y=mx+b m=-9(7,0) the equation of the line is y=

OpenStudy (anonymous):

I came up -63??

OpenStudy (anonymous):

y=-9x+63

OpenStudy (anonymous):

Just plug it in. Y=mx+b so plug in 7 as x 0 as y and -9 as m to find b. So 0=-9(7)+b so 0=-63+b so b=63. so Y=-9x+63